Claudia has sold her car and received approval from the garage owner to re-lease her downtown reserved parking spot for the next four months so she can make some extra money. The rental fee is $200 per month, and she expects to charge $18 per day. Transportation in a car pool will cost her $6 per day. If there are a maximum of 20 work days per month for re-leasing the spot, determine the following:

a. Total cost and revenue relations

b. Breakeven quantity per month

c. Amount of money she will make (or lose) if the number of re-leased days per month over the four-month period are 18, 12, 17, and 20


1
Expert's answer
2021-10-27T08:30:16-0400

Solution:

a.). Given information:

• Rental cost is $200 per month.

• Parking charges (RC) is $18 per day.

• Transportation cost (TC) in a car pool is $6 per day.

• Working days in a month (WD) is 20 days.

Total cost = Total fixed cost + Variable cost

Total fixed cost = 200 "\times" 4 = $800

Variable costs = $6 per day = (6 "\times"20) = $120 "\times"4 = $480

Total monthly cost = 200 + 6x

Total cost (4 months) = 800 + 24x

X = No. of days

Total cost (4 months) = 800 + 24(20) = 800 + 480 = $1,280

Revenue (monthly) = 18x = 18(20) = $360

Revenue (4 months) = 72x = 72(20) = $1,440

 

b.). Break-even quantity per month = "= \\frac{Fixed\\;costs}{(Revenue\\;per\\;day - Variable\\;costs\\;per\\;day)} \n\n\n\u200b"


Break-even quantity per month = "\frac{200}{(18 - 6)} = \frac{200}{12} = 16.7 \;days"


c.). Profit or loss (18 days) = Revenue – Total costs

= (18 "\times" 18) – (200 + (18 "\times" 6)) = 324 – 308 = $16

Profit = $16

Profit or loss (12 days) = Revenue – Total costs

= (18 "\times" 12) – (200 + (12 "\times" 6)) = 216 – 272 = ($56)

Loss = ($56)

 

Profit or loss (17 days) = Revenue – Total costs

= (18 "\times" 17) – (200 + (17 "\times" 6)) = 306 – 302 = $4

Profit = $4

 

Profit or loss (20 days) = Revenue – Total costs

= (18 "\times" 20) – (200 + (20 "\times" 6)) = 360 – 320 = $40

Profit = $40